OpenStudy (anonymous):

find the diagonal of a square whose sides are of the given measure. Given = 16"

OpenStudy (anonymous):

Please look at this resource: http://openstudy.com/study#/updates/4f3571c8e4b0fc0c1a0c78cd Thanks and choose as best answer

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

Pythagorus says \[a^2+b^2=c^2\] i.e. \[16^2+16^2=d^2\] solve this for \(d\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Give me medal because I help to